# LLM Meeting Decision Trees on Tabular Data

### v1

@@ Abstract
Tabular data have been playing a vital role in diverse real-world fields, including healthcare, finance, etc. With the recent success of Large Language Models (LLMs), early explorations of extending LLMs to the domain of tabular data have been developed. Most of these LLM-based methods typically first serialize tabular data into natural language descriptions, and then tune LLMs or directly infer on these serialized data. However, these methods suffer from two key inherent issues: (i) data perspective: existing data serialization methods lack universal applicability for structured tabular data, and may pose privacy risks through direct textual exposure, and (ii) model perspective: LLM fine-tuning methods struggle with tabular data, and in-context learning scalability is bottle-necked by input length constraints (suitable for few-shot learning). This work explores a novel direction of integrating LLMs into tabular data throughough logical decision tree rules as intermediaries, proposes a decision tree enhancer with LLM-derived rule for tabular prediction, DeLTa. The proposed DeLTa avoids tabular data serialization, and can be applied to full data learning setting without LLM fine-tuning. Specifically, we leverage the reasoning ability of LLMs to redesign an improved rule given a set of decision tree rules. Furthermore, we provide a calibration method for original decision trees via new generated rule by LLM, which approximates the error correction vector to steer the original decision tree predictions in the direction of “errors” reducing. Finally, extensive experiments on diverse tabular benchmarks show that our method achieves state-of-the-art performance.
@@ 1 Introduction
Tabular data, typically organized in a structured table format within a relational database with rows and columns standing for the data samples and heterogeneous features (e.g., categorical and numerical features), is fundamental in various real-world fields, including healthcare hernandez2022synthetic , advertising frosch2010decade , finance assefa2020generating , etc. Given the heterogeneity of features borisov2022deep ; gorishniy2021revisiting , decision tree-based methods chen2016xgboost ; ke2017lightgbm ; prokhorenkova2018catboost were found to be particularly suitable for tabular data grinsztajn2022tree . While deep learning has led to breakthroughs in computer vision he2016deep and natural language processing devlin2018bert , decision tree-based methods still outperform the majority of existing deep tabular methods on tabular prediction tasks such as classification and regression. This superiority of decision tree-based methods is explained by the feature heterogeneity for tabular data, where neural networks struggle to learn the irregular target functions compared to decision tree-based methods grinsztajn2022tree .
@@ 2 Related work
Machine learning for tabular prediction. The development of effective algorithms for predictive modeling on tabular data has been a longstanding research topic. In the early days, decision tree-based methods (e.g. XGBoost chen2016xgboost , CatBoost prokhorenkova2018catboost ) were found to be particularly suitable for tabular data. More recently, inspired by the success of of deep learning in computer vision (CV) he2016deep and natural language processing (NLP) devlin2018bert , numerous methods have been proposed for tabular data to accomplish tabular prediction tasks. These works mainly include MLP-like models wang2021dcn , attention-based architectures song2019autoint ; gorishniy2021revisiting ; somepalli2021saint , and retrieval-augmented architectures ye2025revisiting . Among these works, DCN V2 wang2021dcn is an architecture that consists of an MLP-like module and a feature crossing module; AutoInt song2019autoint leveraged the Transformer architecture to capture inter-column correlations; FT-Transformer gorishniy2021revisiting further enhanced AutoInt’s performance through improved token embeddings; ModernNCA ye2025revisiting makes predictions based on the relationships with neighbors in a learned embedding space. Recently, another line of research has tried to use additional information outside target dataset to enhance tabular data prediction. XTab zhu2023xtab pretrains Transformer on a variety of datasets for cross-table pretraining. TabPFN hollmann2025accurate ; hollmann2023tabpfn , which is pretrained on a large set of synthetic datasets, serves as a foundation model for small to medium-sized tabular data.
@@
Large language models for tabular prediction. Motivated by the impressive success of LLMs brown2020language ; achiam2023gpt , another promising study has attempted to harness the rich prior knowledge encapsulated by LLMs in tabular prediction tasks. LIFT dinh2022lift investigated the performance of fine-tuned GPT-3 models brown2020language on tabular data, revealing that the performance of fine-tuned LLMs was roughly on par with traditional solutions. Extending this line of research, TabLLM hegselmann2023tabllm employed T0 sanh2022multitask as the base LLM and demonstrated competitive performance of fine-tuned LLMs in very few-shot scenarios. Additionally, recent efforts have focused on pre-training LLMs over diverse tabular datasets from different domains. Among these, TP-BERTa yan2024making introduces a tabular-specific tokenization scheme, enabling a single pre-trained language model to generalize across multiple tabular datasets after further fine-tuning. GTL wen2024supervised further promotes comprehensive instruction-following capabilities for both zero-shot and in-context learning with a limited number of examples. Despite their promising results, these approaches typically require training or fine-tuning LLMs. Instead of directly fine-tuning LLMs, P2T nam2024tabular leverages unlabeled data correlated with target data expressed in natural language form and prompts LLMs for few-shot semi-supervised learning. Recently, Summary manikandan2023language proposed a boosting framework that treats LLMs as weak learners for tabular prediction, particularly for tasks involving small numbers of data points. FeatLLM han2024large prompts LLMs with serialized training examples to generate new features for few-shot classification tasks. SERSAL yan2025small explores to use synergy learning with FT-Transformer to enhance the noisy annotations generated by LLMs for binary classification tasks in unsupervised manner. This method also needs to fine-tune the LLMs. Going beyond tabular prediction, Nam et al. nam2024optimized leverages LLMs to optimize the feature generator with tree rules in the field of feature engineering zhang2023openfe , which we do not consider as a close related work to ours as the primary goal, motivation, and methodology are different.
@@ 4.1 LLM-based decision tree rules refinement
where $F(x)$ denotes the label prediction produced by the Random Forest, $K$ is the number of experts, $\mathcal{D}_{train}^{k}$ denotes the subset, $r_{k}$ denotes the rule for expert $f_{k}$ derived from Eq. 1, and =$\mathcal{R}=\{r_{k}\}_{k=1}^{K}$ denotes a source decision tree rule set. In terms of decision tree rule, each rule represents a logical comparison between features and thresholds displayed in semantically rich syntax, which thus can be used to partition the feature space into disjoint regions (i.e., leaf nodes) with high interpretability. A well understanding of these rules could generate a better feature space partitioning method, thereby promoting statistical coherence among samples assigned to the same leaf node. Although Random Forest ensembles the outputs from all rules, the inherent relationships and interactions among rules in $\mathcal{R}$ are ignored. With the development of $K$, analyzing these independent rules is gradually becoming more and more difficult, let alone utilizing these rules to partition feature space. To this end, we propose to leverage LLMs to analyze and summarize the rule set $\mathcal{R}$ into a refined rule due to the powerful logical reasoning ability of LLM.
@@
where $\oplus$ is the concatenation of each prompt, and this process is achieved through querying the LLM, without requiring any fine-tuning. Notably, our approach does not require additional domain-specific expertise or semantic information, such as explicit feature names or detailed task background knowledge. This makes it applicable to scenarios where such information is unavailable or incomplete. Instead of relying on the intrinsic domain knowledge of tabular tasks, such as those in healthcare or finance, we frame the problem purely as an domain knowledge-agnostic machine learning task. Our method leverages LLMs to reason decision tree rule set $\mathcal{R}$ into an rule $r^{*}$, rather than directly inferring from serialized tabular data. To verify that LLM could generate a better rule, where samples grouped within the same leaf node exhibit greater statistical similarity, we compute the intra-node sample distance over all leaf nodes partitioned by $r^{*}$, and observe that the distance of $r^{*}$ is lower than original rule $r\in\mathcal{R}$, as illustrated in Fig. 2. Next, we will elaborate on how to leverage the new rule $r^{*}$ generated by LLM to enhance tabular predictions. [Uncaptioned image] Figure 2: Average intra-node distance comparison.
@@ 4.2 Refined rule-guided decision tree error correction
Definition 1. Prediction “errors” increasing and reducing. For a given training sample $x$, we approximate the "error" using the gradient $\nabla_{F(x)}\mathcal{L}(F(x),y)$, where $\mathcal{L}$ is the loss function discussed in Section 3. Intuitively, shifting $F(x)$ in the direction of this gradient (“Errors” Increasing) would likely worsen the model’s performance, analogous to how traditional gradients indicate a direction in parameter space that increases the loss ruder2016overview . Therefore, it is natural that we should calibrate $F(x)$ in the direction of negative gradient $-\nabla_{F(x)}\mathcal{L}(F(x),y)$ (“Errors” Reducing).
@@
where $\mathbb{I}(\cdot)$ is an indicator function determining if $x$ of interest would be assigned to leaf node $L_{l}(r^{*})$ based on rule $r^{*}$, $\eta\in\mathbb{R}_{+}$ is the hyperparameter similar to learning rate controlling the step size, and $\Delta_{x}$ approximates $-\eta*\nabla_{F(x)}\mathcal{L}(F(x),y)$. We now explain the proposed error correction strategy by giving a closer look at Eq. 6.
@@
Let $\mathbb{E}\left[\mathcal{L}(F(x),y)\right]$ denote the expected loss, where $F(x)=\frac{1}{K}\sum_{k}^{K}f_{k}(x|\mathcal{D}_{train}^{k},r_{k})$ and each $f_{k}$ corresponds to a decision tree rule $r_{k}$ from the expert-derived rule set =$\mathcal{R}=\{r_{k}\}_{k=1}^{K}$. Given a prompt $p$ that contains $\mathcal{R}$, the expected loss $\mathbb{E}\left[\mathcal{L}(F(x),y)\right]$ could be decreased by querying the LLM with $p$ to generate a refined rule $r^{*}$, which enables approximation of a sample-specific error correction vector $\Delta_{x}$ to guide prediction $F(x)$ in the direction of “errors” reducing.
@@ 5 Experiments
Baselines and implementation details. We conduct a comparative analysis between DeLTa and other prominent methods in the field of tabular prediction. Specifically, we include LLM-based methods such as TabLLM hegselmann2023tabllm , LIFT dinh2022lift , TP-BERTa yan2024making , GTL wen2024supervised , P2T nam2024tabular , FeatLLM han2024large . For LIFT, we employ two versions, i.e., fine-tuning LLMs (LIFT) and in-context learning (LIFT-ICL) versions, as provided in their original paper. In addition, we compare DeLTa with non LLM-based methods, including KNN cover1967nearest , CART loh2011classification , MLP lecun2015deep , Random Forest breiman2001random , XGBoost chen2016xgboost , CatBoost prokhorenkova2018catboost , FT-Transformer gorishniy2021revisiting , TabPFN hollmann2025accurate , ModernNCA ye2025revisiting . More details about the baseline models can be found in Appendix A.2. Implementation details of DeLTa including hyper-parameters are provided in Appendix A.3. For LLM usage, DeLTa adopts GPT-4o as its LLM backbone, yet is designed to be agnostic to the choice of LLMs (see Appendix A.6 for more results). To prevent LLMs from generating task-irrelevant output, we query LLM 10 times by default and use the average. In replicating baselines, GPT-4o is utilized for in-context learning. For baselines require LLM fine-tuning, we employ the LLMs used in their original paper: TabLLM uses T0 sanh2022multitask ; we use the API provided by OpenAI to perform black-box GPT-3.5 fine-tuning for LIFT, where the fine-tuning method is suggested by the original paper; TP-BERTa employs RoBERTa liu2019roberta ; GTL uses the 13B version of LLaMA 2 touvron2023llama . The experiments are run on NVIDIA A100-PCIE-40GB GPU. To reduce the effect of randomness, the reported performance is averaged over 5 independent runs.
5.1 Main Results

### v2

@@ Abstract
Tabular data have been playing a vital role in diverse real-world fields, including healthcare, finance, etc. With the recent success of Large Language Models (LLMs), early explorations of extending LLMs to the domain of tabular data have been developed. Most of these LLM-based methods typically first serialize tabular data into natural language descriptions, and then tune LLMs or directly infer on these serialized data. However, these methods suffer from two key inherent issues: (i) data perspective: existing data serialization methods lack universal applicability for structured tabular data, and may pose privacy risks through direct textual exposure, and (ii) model perspective: LLM fine-tuning methods struggle with tabular data, and in-context learning scalability is bottle-necked by input length constraints (suitable for few-shot learning). This work explores a novel direction of integrating LLMs into tabular data through logical decision tree rules as intermediaries, proposing a decision tree enhancer with LLM-derived rule for tabular prediction, DeLTa. The proposed DeLTa avoids tabular data serialization, and can be applied to full data learning setting without LLM fine-tuning. Specifically, we leverage the reasoning ability of LLMs to redesign an improved rule given a set of decision tree rules. Furthermore, we provide a calibration method for original decision trees via new generated rule by LLM, which approximates the error correction vector to steer the original decision tree predictions in the direction of “errors” reducing. Finally, extensive experiments on diverse tabular benchmarks show that our method achieves state-of-the-art performance. The source code is available at https://github.com/HangtingYe/DeLTa.
@@ 1 Introduction
Tabular data, typically organized in a structured table format within a relational database with rows and columns standing for the data samples and heterogeneous features (e.g., categorical and numerical features), is fundamental in various real-world fields, including healthcare hernandez2022synthetic, advertising frosch2010decade, finance assefa2020generating, etc. Given the heterogeneity of features borisov2022deep; gorishniy2021revisiting; ye2024ptarl; ye2025drl, decision tree-based methods chen2016xgboost; ke2017lightgbm; prokhorenkova2018catboost were found to be particularly suitable for tabular data grinsztajn2022tree; fan2020autofs; fan2021interactive. While deep learning has led to breakthroughs in computer vision he2016deep and natural language processing devlin2018bert, decision tree-based methods still outperform the majority of existing deep tabular methods on tabular prediction tasks such as classification and regression. This superiority of decision tree-based methods is explained by the feature heterogeneity for tabular data, where neural networks struggle to learn the irregular target functions compared to decision tree-based methods grinsztajn2022tree.
@@ 2 Related work
Machine learning for tabular prediction. The development of effective algorithms for predictive modeling on tabular data has been a longstanding research topic. In the early days, decision tree-based methods (e.g. XGBoost chen2016xgboost, CatBoost prokhorenkova2018catboost) were found to be particularly suitable for tabular data. More recently, inspired by the success of of deep learning in computer vision (CV) he2016deep and natural language processing (NLP) devlin2018bert, numerous methods have been proposed for tabular data to accomplish tabular prediction tasks. These works mainly include MLP-like models wang2021dcn; gorishniy2021revisiting; klambauer2017self, attention-based architectures song2019autoint; gorishniy2021revisiting; somepalli2021saint; arik2021tabnet; huang2020tabtransformer, and retrieval-augmented architectures ye2025revisiting; somepalli2021saint; kossen2021self; gorishniy2023tabr. Among these works, DCN V2 wang2021dcn is an architecture that consists of an MLP-like module and a feature crossing module; AutoInt song2019autoint leveraged the Transformer architecture to capture inter-column correlations; FT-Transformer gorishniy2021revisiting further enhanced AutoInt’s performance through improved token embeddings; ModernNCA ye2025revisiting makes predictions based on the relationships with neighbors in a learned embedding space. Recently, another line of research has tried to use additional information outside target dataset to enhance tabular data prediction. XTab zhu2023xtab pretrains Transformer on a variety of datasets for cross-table pretraining. TabPFN hollmann2025accurate; hollmann2023tabpfn, which is pretrained on a large set of synthetic datasets, serves as a foundation model for small to medium-sized tabular data.
@@
Large language models for tabular prediction. Motivated by the impressive success of LLMs brown2020language; achiam2023gpt; guo2024ds; guo2025optimizing, another promising study has attempted to harness the rich prior knowledge encapsulated by LLMs in tabular prediction tasks. LIFT dinh2022lift investigated the performance of fine-tuned GPT-3 models brown2020language on tabular data, revealing that the performance of fine-tuned LLMs was roughly on par with traditional solutions. Extending this line of research, TabLLM hegselmann2023tabllm employed T0 sanh2022multitask as the base LLM and demonstrated competitive performance of fine-tuned LLMs in very few-shot scenarios. Additionally, recent efforts have focused on pre-training LLMs over diverse tabular datasets from different domains. Among these, TP-BERTa yan2024making introduces a tabular-specific tokenization scheme, enabling a single pre-trained language model to generalize across multiple tabular datasets after further fine-tuning. GTL wen2024supervised further promotes comprehensive instruction-following capabilities for both zero-shot and in-context learning with a limited number of examples. Despite their promising results, these approaches typically require training or fine-tuning LLMs. Instead of directly fine-tuning LLMs, P2T nam2024tabular leverages unlabeled data correlated with target data expressed in natural language form and prompts LLMs for few-shot semi-supervised learning. Recently, Summary manikandan2023language proposed a boosting framework that treats LLMs as weak learners for tabular prediction, particularly for tasks involving small numbers of data points. FeatLLM han2024large prompts LLMs with serialized training examples to generate new features for few-shot classification tasks. SERSAL yan2025small explores to use synergy learning with FT-Transformer to enhance the noisy annotations generated by LLMs for binary classification tasks in unsupervised manner. This method also needs to fine-tune the LLMs. Going beyond tabular prediction, Nam et al. nam2024optimized leverages LLMs to optimize the feature generator with tree rules in the field of feature engineering zhang2023openfe, which we do not consider as a close related work to ours as the primary goal, motivation, and methodology are different.
@@ 4.1 LLM-based decision tree rules refinement
where $F(x)$ denotes the label prediction produced by the Random Forest, $K$ is the number of experts, $\mathcal{D}_{train}^{k}$ denotes the subset, $r_{k}$ denotes the rule for expert $f_{k}$ derived from Eq. 1, and $\mathcal{R}=\{r_{k}\}_{k=1}^{K}$ denotes a source decision tree rule set. In terms of decision tree rule, each rule represents a logical comparison between features and thresholds displayed in semantically rich syntax, which thus can be used to partition the feature space into disjoint regions (i.e., leaf nodes) with high interpretability. A well understanding of these rules could generate a better feature space partitioning method, thereby promoting statistical coherence among samples assigned to the same leaf node. Although Random Forest ensembles the outputs from all rules, the inherent relationships and interactions among rules in $\mathcal{R}$ are ignored. With the development of $K$, analyzing these independent rules is gradually becoming more and more difficult, let alone utilizing these rules to partition feature space. To this end, we propose to leverage LLMs to analyze and summarize the rule set $\mathcal{R}$ into a refined rule due to the powerful logical reasoning ability of LLM.
@@
where $\oplus$ is the concatenation of each prompt, and this process is achieved through querying the LLM, without requiring any fine-tuning. Notably, our approach does not require additional domain-specific expertise or semantic information, such as explicit feature names or detailed task background knowledge. This makes it applicable to scenarios where such information is unavailable or incomplete. Instead of relying on the intrinsic domain knowledge of tabular tasks, such as those in healthcare or finance, we frame the problem purely as a domain knowledge-agnostic machine learning task. Our method leverages LLMs to reason decision tree rule set $\mathcal{R}$ into an rule $r^{*}$, rather than directly inferring from serialized tabular data. To verify that LLM could generate a better rule, where samples grouped within the same leaf node exhibit greater statistical similarity, we compute the intra-node sample distance over all leaf nodes partitioned by $r^{*}$, and observe that the distance of $r^{*}$ is lower than original rule $r\in\mathcal{R}$, as illustrated in Fig. 2. Next, we will elaborate on how to leverage the new rule $r^{*}$ generated by LLM to enhance tabular predictions. [Uncaptioned image] Figure 2: Average intra-node distance comparison.
@@ 4.2 Refined rule-guided decision tree error correction
Definition 1. Prediction “errors” increasing and reducing. For a given training sample $x$, we approximate the "error" using the gradient $\nabla_{F(x)}\mathcal{L}(F(x),y)$, where $\mathcal{L}$ is the loss function discussed in Section 3. Intuitively, shifting $F(x)$ in the direction of this gradient (“Errors” Increasing) would likely worsen the model’s performance, analogous to how traditional gradients indicate a direction in parameter space that increases the loss ruder2016overview. Therefore, it is natural that we should calibrate $F(x)$ in the direction of negative gradient $\nabla_{F(x)}\mathcal{L}(F(x),y)$ (“Errors” Reducing).
@@
where $\mathbb{I}(\cdot)$ is an indicator function determining if $x$ of interest would be assigned to leaf node $L_{l}(r^{*})$ based on rule $r^{*}$, $\eta\in\mathbb{R}_{+}$ is the hyperparameter similar to learning rate controlling the step size, and $\Delta_{x}$ approximates $\eta*\nabla_{F(x)}\mathcal{L}(F(x),y)$. We now explain the proposed error correction strategy by giving a closer look at Eq. 6.
@@
Let $\mathbb{E}\left[\mathcal{L}(F(x),y)\right]$ denote the expected loss, where $F(x)=\frac{1}{K}\sum_{k}^{K}f_{k}(x|\mathcal{D}_{train}^{k},r_{k})$ and each $f_{k}$ corresponds to a decision tree rule $r_{k}$ from the expert-derived rule set $\mathcal{R}=\{r_{k}\}_{k=1}^{K}$. Given a prompt $p$ that contains $\mathcal{R}$, the expected loss $\mathbb{E}\left[\mathcal{L}(F(x),y)\right]$ could be decreased by querying the LLM with $p$ to generate a refined rule $r^{*}$, which enables approximation of a sample-specific error correction vector $\Delta_{x}$ to guide prediction $F(x)$ in the direction of “errors” reducing.
@@ 5 Experiments
Baselines and implementation details. We conduct a comparative analysis between DeLTa and other prominent methods in the field of tabular prediction. Specifically, we include LLM-based methods such as TabLLM hegselmann2023tabllm, LIFT dinh2022lift, TP-BERTa yan2024making, GTL wen2024supervised, P2T nam2024tabular, FeatLLM han2024large. For LIFT, we employ two versions, i.e., fine-tuning LLMs (LIFT) and in-context learning (LIFT-ICL) versions, as provided in their original paper. In addition, we compare DeLTa with non LLM-based methods, including KNN cover1967nearest, CART loh2011classification, MLP lecun2015deep, Random Forest breiman2001random, XGBoost chen2016xgboost, CatBoost prokhorenkova2018catboost, FT-Transformer gorishniy2021revisiting, TabPFN hollmann2025accurate, ModernNCA ye2025revisiting. More details about the baseline models can be found in Appendix A.2. Implementation details of DeLTa including hyper-parameters are provided in Appendix A.3. For LLM usage, DeLTa adopts GPT-4o as its LLM backbone, yet is designed to be agnostic to the choice of LLMs (see Appendix A.6 for more results). To prevent LLMs from generating task-irrelevant output, we query LLM 10 times by default and use the average. In replicating baselines, GPT-4o is utilized for in-context learning. For baselines require LLM fine-tuning, we employ the LLMs used in their original paper: TabLLM uses T0 sanh2022multitask; we use the API provided by OpenAI to perform black-box GPT-3.5 fine-tuning for LIFT, where the fine-tuning method is suggested by the original paper; TP-BERTa employs RoBERTa liu2019roberta; GTL uses the 13B version of LLaMA 2 touvron2023llama. The experiments are run on NVIDIA A100-PCIE-40GB GPU.